OpenStudy (anonymous):

Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 2, – 16), and x intercepts at x = -6 and x = 2. (Do not include the negative sign in your answer.) y = x2 + 4x − ___

OpenStudy (amistre64):

plug in the x values and see where that leads you

OpenStudy (amistre64):

y = x2 + 4x − n fill in the xs to get: y = s - n ; and you want y to equal 0 so; y-s = -n s-y = n

OpenStudy (amistre64):

or something along those lines